Suppose you are trying to save money for a vacation in two years. Bank A is offering savings account with annual interest ratPlease explain how they got 2.25%
Answer #1

Let P be the deposit amount

Let interest rate offered by Bank B is I%, then

F = P *(1+I)^t

F = P *(1+I)^2

Future worth of deposit in Bank A = P *(1+0.02) *(1+0.025) = P *(1.02) *(1.025) = P * 1.0455

As per condition given, both these deposits have to be equivalent, so

P * 1.0455 = P *(1+I)^2

(1+I)^2 = 1.0455

1+i = 1.0455 ^ (1/2) = 1.0224969

I = 1.0224969 - 1 = 0.0224969

I = 0.0225 = 2.25%

Therefore correct answer is option D